# Ruelle-Pollicott resonances of diffusive U(1)-invariant qubit circuits

### v1

@@ Appendix C Ruelle-Pollicott resonances through weak dephasing
In this section we demonstrate the weak Lindbladian dissipation-based method, proposed in Ref. Mori (2024), on circuits without any conservation, and briefly show it gives the same results as the truncated propagator method. Additionally, we show that in circuits we studied, the convergence of the Lindblad method can sometimes be rather tricky.
The idea is Mori (2024) to introduce some kind of Lindbladian dissipation of strength $\gamma$, obtaining the open system version of the Heisenberg propagator $\mathcal{U}_{\gamma}$ (which, in this case, is the exponent of the Lindbladian). The leading RP resonances $\lambda$ then correspond to the leading eigenvalues of the Heisenberg propagator $\lambda_{\mathcal{U}_{\gamma}}$ in the zero noise $\gamma\to 0^{+}$ limit and in the thermodynamic limit, i.e.,
$\lambda=\lim_{\gamma\to 0^{+}}\lim_{N\to\infty}\lambda_{\mathcal{U}(\gamma)},$ (34)
where $N$ is the system size. The order of limits is important, the opposite order always gives 00.
@@
$\mathcal{D}(A)$ :-$\coloneq\gamma\sum_{j=1}^{N}\left(\sigma^{z}_{j}A\sigma^{z}_{j}-A\right),:$ (35)
$\mathcal{U}(A)$ :-$\coloneq{\rm e}^{\mathcal{D}}(U^{\dagger}AU),:$
@@
One can calculate the leading eigenvalue of $\mathcal{U}_{\gamma}$ with an iterative solver (similar to what we did with the truncated propagator, described in Appendix B.1) for different $N$ and $\gamma$. Extrapolation to zero noise and the thermodynamic limit must be done carefully Mori (2024). Namely, if one observes the curve $|\lambda_{\mathcal{U}}(\gamma)|$ for fixed $N$, we see that it “breaks” (i.e., abruptly changes its slope) at some small $\gamma$. This change has a physical origin: for small $\gamma$ the scattering length becomes larger than the system size and one starts to be influenced by finite-size (boundary) effects. Therefore, in order to obtain the correct decay rate of the bulk physics from a finite-$N$ data one has to extrapolate the curve $|\lambda_{\mathcal{U}}(\gamma)|$ from finite (large) $\gamma$ down to $\gamma=0$.
@@
An example of such behavior is seen in numerical data for one realization of a brickwork circuit (with a rather tricky instance of $V$) shown in Fig. 10a. Instead of $\lambda$ we plot $\nu_{\mathcal{U}}(\gamma)=-\log\left\lvert\lambda_{\mathcal{U}}(\gamma)\right\rvert$, like in Ref. Mori (2024). We observe $\nu_{\mathcal{U}}(\gamma)$ is independent of the system size $N$ for big enough $\gamma\gtrsim 0.15$, which matches the predictions of Ref. Mori (2024). From data at $\gamma>0.15$, we can extrapolate (black dashed curve) to obtain the prediction for the leading RP resonance $\left\lvert\lambda\right\rvert\approx{\rm e}^{-0.403}\approx 0.668$. Comparing this to the value obtained by the truncated propagator (Fig. 10a.i) and to decay of correlation functions (Fig. 10b), we see it is incorrect. The long time decay is actually governed by $\left\lvert\lambda\right\rvert\approx{\rm e}^{-0.181}\approx 0.835$, which the truncated propagator predicts correctly (in this case, this is the maximum $\lambda_{1}(k)$ over all $k$, which occurs at $k=0$).
@@
If we study $\nu_{\mathcal{U}}(\gamma)$ more carefully though, we can get the correct RP resonance from the Lindbladian approach too. Calculating $\nu_{\mathcal{U}}(\gamma)$ for smaller $\gamma$, we in fact observe another breaking of the curve. This time, the curves for different $N$ do not overlap, but we can still extrapolate for each $N$ separately and then try to do $N\to\infty$. Such a procedure is also shown in Fig. 10a (colored dotted curves), with the convergence with $N$ shown in the inset (a.i). While the procedure does not yet converge for accessible $N$, it is in the vicinity of what appears to be the correct RP resonance as obtained from the truncated propagator. Additionally, in Fig. 10 we see that the early time behavior of the magnetization correlation function seems to be governed by $\left\lvert\lambda\right\rvert\approx 0.668$. The initial breaking of the $\nu_{\mathcal{U}}(\gamma)$ therefore might not be spurious, but rather an indication of a subleading resonance.
@@
We must emphasize that the showcased behavior is not generic. Double breaking is not always observed; sometimes the correct RP resonance is given by the first breaking, sometimes not even by the second. Additionally, we do not necessarily see a 2-step relaxation in generic correlation functions. All in all, this example showcases that, in our experience, the Lindbladian method is less robust. Taking the double limit involves making extrapolations that are sometimes hard to make. Furthermore, if we would want to treat systems with a conservation law, like magnetization, the method is much more cumbersome, and it is not clear how to do the momentum resolution.
(a)                                                                  Refer to caption (b)
Figure 11: Finite-size plateaus in the autocorrelation function of $A=\sum_{j}\sigma^{z}_{j}\sigma^{z}_{j+1}$. Shown are numerical results for a circuit with a random 3-site magnetization-conserving gate (Sec. III) and different $N$. Subfigure (a) shows time dependence, while (b) shows the values of plateaus in (a). The red dashed line are the leading asymptotics (41).
@@ Appendix D Plateaus in correlation functions due to the powers of the conserved quantity
Powers of a conserved quantity are also conserved quantities. When studying physical properties of many-body systems they can be often ignored because a power of a local operator is a non-local operator. In finite systems with size $N$, however, their impact can be non-negligible. Namely, local observables can have a finite overlap with some power of a conserved quantity and thus cause its correlation function to saturate to a higher value than the expected fluctuations (see Appendix B.2). Although this mechanism is simple, we explicitly demonstrate it in the case of magnetization conservation, since one needs to be aware of it when interpreting finite-size numerics.
@@
$t\to\infty]{}\left\lvert c_{Q}\right\rvert^{2}\left\langle Q,Q\right\rangle=\frac{\left\langle Q,A\right\rangle^{2}}{\left\langle Q,Q\right\rangle}.$ (37)
@@
Let’s now consider the correlation function of a $2$-site extensive observable
@@
evolved by a magnetization-conserving circuit. $A$ has zero overlap with the conserved $M=\sum_{j}\sigma^{z}_{j}$, so its correlation function should decay to 00 in the thermodynamic limit $N\to\infty$. While this is true, it also has a non-zero overlap with powers of $M$ and its correlation function will therefore exhibit a finite-size plateau.
Although $A$ overlaps with all even powers of magnetization, the leading order of the plateau in $1/N$ will be caused by $M^{2}$. For $Q$ one can now take $M^{2}$ made traceless (i.e., with the identity projected out), $Q=M^{2}-N\mathbbm{1}$. A short calculation gives
@@
The correlation functions of $A$ in circuits with a random 3-site gate (i.e., the one described in Sec. III) and different $N$ are shown in Fig. 11a. The plateaus are shown in Fig. 11b and match the predicted $\sim 1/N$ scaling at large $N$.
@@ Appendix E Stretched-exponential decay of local correlation functions
Recently McCulloch et al. (2025), it was argued that all local correlation functions unrelated to transport decay as stretched exponentials in diffusive one-dimensional systems, that is as $\sim{\rm e}^{-Bt^{\alpha}}$ for some $\nu$ and $\alpha$. For translationally-invariant Floquet systems, Ref. McCulloch et al. (2025) predicts $\alpha=2/3$ and numerically confirms the prediction for magnetization-conserving brickwall circuits.
We check this in the $3-$site generalized brickwall circuit (see Sec. III) by analyzing the local correlation function of $x$ magnetization, which is not expected to be described by an effective (hydrodynamic) theory of transport McCulloch et al. (2025); Matthies et al. (2024). The numerics are shown in Fig. 12, and we average over space to obtain better statistics. The average is done on the absolute value squared, since averaging just $\left\langle\sigma^{x}(i,t)\sigma^{x}(0,0)\right\rangle$ would yield the extensive $x$ magnetization correlation function. While the numerics shown suggest the decay is better described by a stretched exponential with $\alpha=2/3$ than an exponential, fitting the power $\alpha$ has a big uncertainty (we obtain $\alpha\sim 0.4-0.8$ for different models with fitting uncertainties at the order of $\sim 0.1$).
As argued in Sec. IV.2, stretched-exponential decays are governed by RP continuums with closing gap. An interesting open question is what happens for correlation functions of extensive observables, which are essentially the Fourier transform of local observables. Our numerics seem to indicate exponential decay, at least for big $k$ (see Sec. III.2). Asymptotics for correlation functions for $k=0$ are less clear numerically, the extensive $x$ magnetization does not seem to be in the asymptotic regime for accessible time scales and additionally heavily oscillates.
Figure 12: Local correlation function of $x$ magnetization. The correlation functions are calculated in a circuit with a random 3-site magnetization-conserving gate (Sec. III) with $N=30$ sites and averaged over space. Dashed lines show a stretched exponential ${\rm e}^{-Bt^{2/3}}$ fit.
@@ Appendix F Local spin current
In this section, we provide additional details about the diffusive behavior of the local spin current discussed in Sec. IV.2.1. The spin current $j$ is defined by the discrete continuity equation, Eq. (21). The diffusive predictions come from the diffusive local correlation function of magnetization, Eq. (15), and the continuity equation. For clarity, we use the continuous space-time continuity equation
@@
Similarly, as in Sec. III.1, the correlation function of extensive current with good quasi-momentum $J_{k}:=\sum_{l}{\rm e}^{-{\rm i}lk/3}j_{l}$ can now be studied. We obtain
@@
The decay is, therefore, governed by an RP resonance with the same $k$ dependence as the one governing the decay of magnetization [cf. Eq. (16)], with the prefactor $k^{2}$ interpreted as an expansion coefficient.

### v2

@@ Appendix C Ruelle-Pollicott resonances through weak dephasing
In this section, we demonstrate the weak Lindbladian dissipation-based method, proposed in Ref. [29], on circuits without any conservation, and briefly show it gives the same results as the truncated propagator method. Additionally, we show that in the circuits we studied, the convergence of the Lindblad method can sometimes be rather tricky.
The idea is [29] to introduce some kind of Lindbladian dissipation of strength $\gamma$, obtaining the open system version of the Heisenberg propagator $\mathcal{U}_{\gamma}$ (which, in this case, is the exponent of the Lindbladian). The leading RP resonances $\lambda$ then correspond to the leading eigenvalues of the Heisenberg propagator $\lambda_{\mathcal{U}_{\gamma}}$ in the zero noise $\gamma\to 0^{+}$ limit and in the thermodynamic limit, i.e.,
$\lambda=\lim_{\gamma\to 0^{+}}\lim_{N\to\infty}\lambda_{\mathcal{U}_{\gamma}},$ (34)
where $N$ is the system size. The order of limits is important, the opposite order always gives 0 gap.
@@
$\mathcal{D}_{\gamma}(A)$ $\coloneq\gamma\sum_{j=1}^{N}\left(\sigma^{z}_{j}A\sigma^{z}_{j}-A\right),$ (35)
$\mathcal{U}_{\gamma}(A)$ $\coloneq{\rm e}^{\mathcal{D}_{\gamma}}(U^{\dagger}AU),$
@@
(a)                    Refer to caption (b)                  Refer to caption (c)
Figure 10: Extraction of the leading RP resonance by a weak-dephasing Lindbladian [see Eq. (35)], an analogous figure to Fig. 2. (a) The leading Liouvillian eigenvalue for the dissipative brickwall quantum circuit with a Haar-random gate $V$. Dashed and dotted lines denote extrapolation to zero noise. (b) The convergence of the second breaking extrapolation [dotted lines in (a)] compared to the first breaking extrapolation [dashed lines (a)] and the RP resonance estimated from $r=12$ truncated propagator (the biggest gap is at $k=0$). (c) Infinite temperature autocorrelation function of extensive $k=0$ observables, defined in Eq. (4). The local densities are either $\sigma^{z}$ (i.e., $z$ magnetization) or chosen randomly according to the Gaussian unitary ensemble (GUE) [27] with support $r=2$. The dashed and dotted lines denote the decays $\propto\left\lvert\lambda\right\rvert^{t}$ governed by the first and second breaking predictions, respectively. The numerics are done in a finite circuit with $N=32$ sites, the $\sigma^{z}$ correlation function is multiplied by 99 for better presentation.
@@
One can calculate the leading eigenvalue of $\mathcal{U}_{\gamma}$ with an iterative solver (similar to what we did for the truncated propagator, described in Appendix B.1) for different $N$ and $\gamma$. Extrapolation to zero noise and the thermodynamic limit must be done carefully [29]. Namely, if one observes the curve $|\lambda_{\mathcal{U}_{\gamma}}|$ for fixed $N$, we see that it “breaks” (i.e., abruptly changes its slope) at some small $\gamma$. This change has a physical origin: for small $\gamma$ the scattering length becomes larger than the system size and one starts to be influenced by finite-size (boundary) effects. Therefore, in order to obtain the correct decay rate of the bulk physics from finite-$N$ data one has to extrapolate the curve $|\lambda_{\mathcal{U}_{\gamma}}|$ from finite (large) $\gamma$ down to $\gamma=0$.
@@
An example of such behavior is seen in numerical data for one realization of a brickwall circuit (with a rather tricky instance of $V$) shown in Fig. 10a. Instead of $\lambda$ we plot $\nu_{\mathcal{U}_{\gamma}}=-\log\left\lvert\lambda_{\mathcal{U}_{\gamma}}\right\rvert$, like in Ref. [29]. We observe $\nu_{\mathcal{U}_{\gamma}}$ is independent of the system size $N$ for big enough $\gamma\gtrsim 0.15$. From data at $\gamma>0.15$, we can extrapolate (black dashed curve) to obtain the prediction for the leading RP resonance $\left\lvert\lambda\right\rvert\approx{\rm e}^{-0.403}\approx 0.668$. Comparing this to the value obtained by the truncated propagator (Fig. 10b) and to decay of correlation functions (Fig. 10c), we see it is incorrect. The long time decay is actually governed by $\left\lvert\lambda\right\rvert\approx{\rm e}^{-0.181}\approx 0.835$, which the truncated propagator predicts correctly (in this case, this is the maximum $\lambda_{1}(k)$ over all $k$, which occurs at $k=0$).
@@
If we study $\nu_{\mathcal{U}_{\gamma}}$ more carefully, we can get the correct RP resonance from the Lindbladian approach too. Calculating $\nu_{\mathcal{U}_{\gamma}}$ for smaller $\gamma$, we in fact observe another breaking of the curve. This time, the curves for different $N$ do not overlap, but we can still extrapolate for each $N$ separately and then try to do $N\to\infty$. Such a procedure is also shown in Fig. 10a (colored dotted curves), with the convergence with $N$ shown in the inset (a.i). While the procedure does not yet converge for accessible $N$, it is in the vicinity of what appears to be the correct RP resonance as obtained from the truncated propagator. Additionally, in Fig. 10 we see that the early time behavior of the magnetization correlation function seems to be governed by $\left\lvert\lambda\right\rvert\approx 0.668$. The initial breaking of the $\nu_{\mathcal{U}_{\gamma}}$ therefore might not be spurious, but rather an indication of a subleading resonance.
@@
We must emphasize that the showcased behavior is not generic. Double breaking is not always observed; sometimes the correct RP resonance is given by the first breaking, sometimes not even by the second. Additionally, we do not necessarily see a 2-step relaxation in generic correlation functions. All in all, this example showcases our experience that the Lindbladian method is less robust. Taking the double limit involves making extrapolations that are sometimes hard to perform. Furthermore, if we would want to treat systems with a conservation law, like magnetization, the method is much more cumbersome, and it is not clear how to do the momentum resolution efficiently.
@@ Appendix D Plateaus in correlation functions due to the powers of the conserved quantity
Powers of a conserved quantity are also conserved quantities. When studying physical properties of many-body systems, they can be often ignored because a power of a local operator is a non-local operator. In finite systems of size $N$, however, their impact can be non-negligible. Namely, local observables can have a finite overlap with some power of a conserved quantity and thus cause its correlation function to saturate to a higher value than the expected fluctuations (see Appendix B.2). Although this mechanism is simple, we explicitly demonstrate it in the case of magnetization conservation, since one needs to be aware of it when interpreting finite-size numerics.
@@
$\xrightarrow[t\to\infty]{}\left\lvert c_{Q}\right\rvert^{2}\left\langle Q,Q\right\rangle=\frac{\left\langle Q,A\right\rangle^{2}}{\left\langle Q,Q\right\rangle}.$ (37)
@@
Let’s now demonstrate this with the correlation function of a $2$-site extensive observable
@@
evolved by a magnetization-conserving circuit. $A$ has zero overlap with the conserved $M=\sum_{j}\sigma^{z}_{j}$, so its correlation function should decay to 0 in the thermodynamic limit $N\to\infty$. While this is true, it has a non-zero overlap with powers of $M$ and its correlation function will therefore exhibit a finite-size plateau.
Although $A$ overlaps with all even powers of magnetization, the leading order of the plateau in $1/N$ will be caused by $M^{2}$. For $Q$ one takes the traceless part of $M^{2}$, $Q=M^{2}-N\mathbbm{1}$. A short calculation gives
@@
(a)                     Refer to caption (b)
Figure 11: Finite-size plateaus in the autocorrelation function of $A=\sum_{j}\sigma^{z}_{j}\sigma^{z}_{j+1}$. Shown are numerical results for a circuit with a random 3-site magnetization-conserving gate (Sec. 3) and different NN. Subfigure (a) shows time dependence, while (b) shows the values of plateaus in (a). The red dashed line are the leading asymptotics (41).
The correlation functions of $A$ in a homogeneous circuit with a random 3-site gate (i.e., the one described in Sec. 3) and different NN are shown in Fig. 11a. The plateaus are shown in Fig. 11b and match the predicted $\sim 1/N$ scaling at large NN.
@@ Appendix E Stretched-exponential decay of local correlation functions
Recently [26], it was argued that all local correlation functions unrelated to transport decay as stretched exponentials in diffusive one-dimensional systems, that is as $\sim{\rm e}^{-Bt^{\alpha}}$ for some $B$ and $\alpha$. For translationally-invariant Floquet systems, Ref. [26] predicts $\alpha=2/3$ and numerically confirms the prediction for magnetization-conserving brickwall circuits.
We check this in the 3-site generalized brickwall circuit (see Sec. 3) by analyzing the local correlation function of $\sigma^{x}$ magnetization, which is not expected to be described by an effective (hydrodynamic) theory of transport [26, 25]. The numerics are shown in Fig. 12, we average over space to obtain better statistics. The average is done on the absolute value squared, since averaging just $\left\langle\sigma^{x}(i,t)\sigma^{x}(0,0)\right\rangle$ would yield the extensive $\sigma^{x}$ magnetization correlation function. While the numerics shown suggest the decay is better described by a stretched exponential with $\alpha=2/3$ than an exponential, fitting the power $\alpha$ has a big uncertainty (we obtain $\alpha\sim 0.4-0.8$ for different models with fitting uncertainties of the order of $\sim 0.1$).
Figure 12: Local correlation function of $\sigma^{x}$ magnetization. The correlation function is calculated in a circuit with a random 3-site magnetization-conserving gate (Sec. 3) with $N=33$ sites and is averaged over space. Dashed lines show a stretched exponential ${\rm e}^{-Bt^{2/3}}$ fit.
As argued in Sec. 4.2, stretched-exponential decays are governed by RP continuums with closing gap. An interesting open question is what happens for correlation functions of extensive observables, which are essentially the Fourier transform of local observables. Our numerics seem to indicate exponential decay, at least for big $k$ (see Sec. 3.2). Asymptotics for correlation functions for $k=0$ are less clear numerically, the extensive $x$ magnetization does not seem to be in the asymptotic regime for accessible time scales and additionally heavily oscillates.
@@ Appendix F Local spin current
In this section, we provide additional details about the diffusive behavior of the local spin current discussed in Sec. 4.2.1. The spin current $j$ is defined by the discrete continuity equation, Eq. (21). The diffusive predictions come from the diffusive local correlation function of magnetization, Eq. (15), and the continuity equation. For clarity, we use the continuous space-time continuity equation
@@
Similarly, as in Sec. 3.1, the correlation function of extensive current with good quasi-momentum $J_{k}:=\sum_{l}{\rm e}^{-{\rm i}lk/3}j_{l}$ can now be studied. We obtain
@@
The decay is, therefore, governed by an RP resonance with the same $k$ dependence as the one governing the decay of magnetization [cf. Eq. (16)], with the prefactor $k^{2}$ perhaps interpreted as an expansion coefficient.